NewMach Inc.

TRiedTested

MayBEERuns

Annual Fixed Costs

115,000

197,000

78,000

Per Hour Operating

140

125

150

Per Hour Labor Cost

18

21

25

Per Hour Repair

2.6

1.1

2.1

Per Hour Software Charge

1.1

0

0

MakeIt4U does contract manufacturing, large batch production under contract to companies that place their brand on the product. Usually these contracts last at a minimum of 4 months. Data on the costs of the three machine options to buy is shown above.

  1. What is the lowest hour usage that the machine from NewMach is the lowest cost option?     
  2. What is the highest hour usage that the machine from NewMach is the lowest cost option?     
  3. What is the average cost per unit (including fixed cost) using the lowest overall cost machine if the hours of usage are estimated to be 5900?    

Please answer all parts!
Answer #1
NewMach Inc. TRiedTested MayBEERuns
Annual Fixed Costs 115000 197000 78000
Per Hour Operating 140 125 150
Per Hour Labor Cost 18 21 25
Per Hour Repair 2.6 1.1 2.1
Per Hour Software Charge 1.1 0 0
Total cost per hour 161.7 147.1 177.1

Machine MayBEERuns NewMachTRiedTested Fixed Cost (FC)78,000115,000 197,000 147.1 Variable Cost (VC)177.1 161.7 Total Cost FC Q * VC Quantity (MayBEERuns NewMach TRiedTested 0 2,403 5,616 115,000 503,500 1,072,671 1,023,178 1,023,178 78,000 503,500 197,000 550,422 Crossover point between MayBEERuns & NewMach (115000 - 78000) / (177.1 - 161.7) 2403 crossover point between NewMach & TRedTested (197000-115000) / (161.7-147.1-5616 Optimal Range of Output (Q) Alternative MayBEERuns NewMach TRiedTested 0 2,403 5,616 2,403 5,616MayBEERuns Ne 1,400,000 1,200,000 1,000,000 tai 600,000 400,000 200,000 0,0 1,0002,000 3,000 4,000 5,000 6,000 7,000 8,000 NUMBER OF HOURS

a. The lowest hour usage that the machine from NewMach is the lowest cost option = 2403 hours

b. The highest hour usage that the machine from NewMach is the lowest cost option = 5616 hours

c. The average cost per unit (including fixed cost) using the lowest overall cost machine if the hours of usage are estimated to be 5900

For this usage, lowest cost option is TRiedTested.

Total cost = 197000+147.1*5900 = $ 1,064,890

Average cost per unit = 1064890/5900 = $ 180.5